if 60cm3 of a gas is heated from 27'-50'C. whats the new volume at constant pressure
1
Expert's answer
2016-02-10T00:01:18-0500
According to Charles Law the value of V/T is constant at the constant pressure. Hence, the initial state
has V1 = 60 cm3 , T1 = 27 oC = 300 K, and the final characteristics are T2 = 50 oC = 273 + 50 K = 323 K, V2 is
unknown. Let’s find V2 using Charles Law:
V1/T1 = V2/T2
V2 = T2×(V1/T1)
The new volume is : V2 = 323 ×(60/300) = 64.6 cm3